OpenStudy (anonymous):

write an equation in the form y=mx+b for the line which passes through (1, -10) and (-7, -2)

OpenStudy (anonymous):

the slop is -8/-6

OpenStudy (anonymous):

u ther

OpenStudy (anonymous):

thank you..

OpenStudy (anonymous):

1) first find the slope, so use the formula y2-y1/x2-x1, so ur slope is -1 2) then use point slope formula y-y1=m(x-x1) 3) choose any coordinate and plug it in 4) the equation would look like y+2=-1(x+7) 5) then the equation is y+2=-1x-7 6) add -2 to both sides to get a final equation of y=-x-9 hope that helps :D